OpenStudy (anonymous):

use factoring to solve the quadratic equation: x(x-8)=-16(x+1)

OpenStudy (anonymous):

x^2+8x+16=0 => (x+4)(x+4)=0 so, x=-4,-4

OpenStudy (anonymous):

so, with these problems, you will always have to put it in ( )( ) form?

OpenStudy (anonymous):

yes, factorise and keep 0 to the rhs

OpenStudy (anonymous):

okay thanks =] so with x(x+22)=11(4x-11) it would be (x+11)(x-2)=0 ?